OpenStudy (anonymous):

Find an equation of the curve through (1,1) the slope of whose tangent line at (x,y) is y^9x^6. Express your answer as y^-8=

OpenStudy (anonymous):

The wording is a bit different but think of it as just a differential equations problem: dy/dx=y^9x^6 and y(1)=1

OpenStudy (anonymous):

So after I separate and integrate each side I should get -1/8y^8=(-1/5x^5)+C right?

OpenStudy (abb0t):

Looks like basic algebra to me.